# Dimensional reduction and adaptation-development-evolution relation in evolved biological systems

### v1

@@ 7 Homeorhesis in Development and Evolution-Development Relationship
A promising extension of dimensional reduction and relationship will be in the area of evolutionary development.
To consider the robustness of the developmental process, Waddington, in the 1950s, schematically represented the developmental process by the motion of a ball along a landscape of branching valleys, as shown in Fig. 4[41]. The ball descends into lower valleys corresponding to stable cell types, while the successive branching of the valleys represents the differentiation process. In the figure, the horizontal axis represents the cell state that gives rise to the cell type, and the depth axis represents a slow developmental time. This picture has been revived in recent years and has received much attention.
Here, the horizontal axis represents the cellular state, possibly represented by the gene expression pattern. However, since there are thousands of genes (or components) in a cell, the cellular state involves many degrees of freedom, as mentioned above, while the horizontal axis represents only one variable. So how can cellular states potentially be represented by only one (or a few) variables? This first question is again a question of dimensional reduction, possibly achieved by evolution.
Since the stable state of each differentiated cell type is at the bottom of the valley, this may fit the attractor picture of a cellular state: For a given value of the depth axis , the cellular state is attracted to the bottom of the valley, where the state falls onto a fixed-point attractor, and no further change occurs. Here, however, the landscape itself is shaped along the depth axis with the developmental course, with which the valleys are successively shaped and deepened in a process known as canalization. Then the second question we need to address is: what does the y-axis represent for the (slower) landscape change and what is the origin of such a slow mode[42]?
Third, each developmental process is robust, i.e., how and when the valleys branch is rater robust. Waddington coined the term homeorhesis to discuss this robustness in the path[41]. This is not the stability of a state; how a path is robust for evolution is the third question that remains to be answered.
@@
In this context, the concept of the developmental hourglass has recently gained much attention, in which different species within the same phylum increase their similarity in the mid-developmental stage. The difference between different species is large in the egg stage, then shrinks in the mid-developmental stage, and then increases again to have diverse adult forms. Because the bottleneck narrows in the mid-stage, it is called the hourglass hypothesis. When it was first proposed[46], it was difficult to determine quantitatively, but recently it has become possible to quantitatively compare gene expression profiles at each stage. The data seem to support the hypothesis [47, 48, 49], while further data analysis is needed to draw clear conclusions. Now, can we understand the developmental hourglass theoretically?
@@
3) Evolution of robustness in the developmental path rather than robustness of the final state.
4) Correlation between developmental and evolutionary changes, rather than correlation between the time course of phenotypic variances due to noise or environmental changes and phenotypic variation due to genetic changes, over developmental time .
To understand this evolutionary developmental relationship, we performed evolutionary simulations of the developmental process [50, 51] Specifically, cells follow the gene expression dynamics governed by the identical gene regulatory network. These cells are placed in one dimension, while some components are assumed to be exchanged with neighboring cells through diffusion interactions. Furthermore, external chemical components (morphogens) are supplied with a spatial gradient, providing input to the gene expression system. Under these conditions, this gene expression response and diffusion provides the dynamics for the gene expression level (protein concentration) in each cell. In the model, the gene expression levels of each cell reach a steady state that depends on the location of the cell, leading to pattern formation[50].
Let us then assume that the fitness is given by the number of final cell types defined by the expression pattern of target genes (specifically four target genes are adopted, so that cell types are possible according to the possibilities of all their on-off patterns.) By using the fitness and mutating the gene regulation networks over generations, the individuals with higher fitness are selected for the next generation[51].
After a few hundred generations, the networks have evolved to reach the fittest. It turns out that during evolution, the expression of a certain gene (protein) slowly changes, controlling the expression of the output genes. (For details, see the original paper[51]).
@@
C) Behind the emergence of the developmental hourglass in the model, there exists the acquisition of genes which govern a slow scheduling process in development and induce developmental robustness. Such a slow mode controls the expression dynamics of other genes, which also facilitates evolution. In fact, these slow expression dynamics were conserved among organisms within a certain phylogenetic range of species sharing the hourglass (point 2).
D) Thus, the developmental hourglass can be related to the robustness and dimensional reduction by a slow mode as discussed in Β§2: The phenotypic diversity inherited from the parents at the egg stage must be reduced at some point, in order for the developmental process to proceed stably (point 3). This leads to the low-dimensional constraint on the developmental path of phenotype (point 1), leading to the decrease in variance at mid-developmental stage, the timing of which is given by this slowly changing gene expression. By presuming that it is correlated with the the the variance over mutants through the developmental time. the developmental bottleneck over species will be expected (point 4)[52].
@@ 8 Collapse of dimensional reduction: Anna Karenina phenomena?
So far we have discussed the achievement of dimensional reduction in a biologically good state. Then, it will be of interest to study if the reduction remains to be valid for cells with a bad, non-growing state which are put into poor environmental conditions. In fact, in bacteria, there generally exists a transition to the dormant state, i,e,, the growth-arrested state, when the nutrient concentration is decreased[53, 54], whereas theoretical models for the transition from an exponential growth state to a quiescent state (which cannot grow) have also been proposed[56, 55]. Will the dimensional reduction collapse with this transition to the dormant state?
@@
If the rate of the second reaction is much faster, the product is formed as soon as the complex is formed, so that the above reaction processes are reduced to single mass action kinetics as , without intermediate complex formation. Then the model is reduced to the cell model mentioned already in Β§2-6. On the other hand, if the rate is not so fast, the intermediate complex remains formed for some time, and the complexes may accumulate depending on the condition. If such accumulation occurs, the abundances of free reactants that are not bound in complexes will decrease, which can hinder the reaction processes.
@@
In fact, an increase in the diversity of chemical components is observed as the cell enters a non-growing quiescent state. This increase in diversity suggests a violation of dimensional reduction in an exponential growth state. In a bad state, the dimensional reduction achieved for a robust state with consistency between macro and micro levels may collapse, leading to diversity in components. Probably, the first sentence of Tolstoyβs Anna Karenina, βHappy families are all alike; every unhappy family is unhappy in its own wayβ, can be applied to cells or biological systems.
@@ 9 Summary
In this paper, we have reviewed recent advances in evolutionary dimension reduction. Noting that the biological system is robust to perturbations while maintaining plasticity for evolution, the phenotypic changes, which can originally be high-dimensional, are constrained to low-dimensional space. The responses to environmental changes, stochastic perturbations, and genetic changes follow dominantly along this low-dimensional manifold, and thus the responses to environmental and genetic changes, fluctuations to noise, and to genetic mutations are strongly correlated, allowing the prediction of phenotypic evolution. The extension of the dimensional reduction to the development-evolution relationship is briefly reviewed. Finally, a possible collapse of the reduction in a dormant (non-growing) cellular state is briefly discussed.

### v2

@@ 7 Homeorhesis in Development and Evolution-Development Relationship
A promising extension of dimensional reduction and - relationship will be in the area of evolutionary development. To consider the robustness of the developmental process, Waddington, in the 1950s, schematically represented the developmental process by the motion of a ball along a landscape of branching valleys, as shown in Fig. 4[51]. The ball falls down into lower valleys that represent stable cell types, while the successive branching of the valleys represents the differentiation process. In the figure, the horizontal axis represents the cell state that gives rise to the cell type, and the depth axis represents a slow developmental time. This picture has been revived in recent years and has received much attention.
Here, the horizontal axis represents the cellular state, possibly represented by the gene expression pattern. However, since there are thousands of genes (or components) in a cell, the cellular state involves many degrees of freedom, as already mentioned, while the horizontal axis represents only one variable. Then, how can cellular states potentially be represented by only one (or a few) variables? This first question is again a question of dimensional reduction, which can possibly be achieved by evolution.
Since the stable state of each differentiated cell type is at the bottom of the valley, this may fit the attractor picture of a cellular state: For a given value of the depth axis , the cellular state is attracted to the bottom of the valley, where the state falls onto a fixed-point attractor, and no further change occurs. Here, however, the landscape itself is shaped along the depth axis with the developmental course, with which the valleys are successively shaped and deepened in a process known as canalization. Then, the second question we need to address is: what does the y-axis represent for the (slower) landscape change and what is the origin of such a slow mode[52]?
Third, developmental process is robust, i.e., how and when the valleys branch is rater robust. Waddington coined the term homeorhesis to discuss this robustness in the path[51]. This is not the stability of a state; how a developmental path is robust is the third question that remains to be answered.
@@
In this context, the concept of the developmental hourglass has recently gained much attention, in which different species within the same phylum increase their similarity in the mid-developmental stage. The difference in phenotypes between different species is large in the egg stage, then shrinks in the mid-developmental stage, and later increases again to have diverse adult forms. Because the bottleneck narrows in the mid-stage, it is called the hourglass hypothesis. When it was first proposed[56], it was difficult to determine quantitatively, but recently it has become possible to quantitatively compare gene expression profiles at each stage. The data seem to support the hypothesis [57, 58, 59, 60], while further data analysis is needed to draw clear conclusions. Now, can we understand the developmental hourglass theoretically?
@@
3) Evolution of robustness of the developmental path rather than robustness of the final state.
4) Correlation between the time course of phenotypic variances due to noise or environmental changes and due to genetic changes, over developmental time , rather than the correlation between the steady-state variances and .
To understand this evolutionary developmental relationship, we performed evolutionary simulations of the developmental process [61, 62] Specifically, cells follow the gene expression dynamics governed by the identical gene regulatory network. These cells are placed in one dimension, while some components are assumed to be exchanged with neighboring cells through diffusive interactions. Furthermore, external chemical components (morphogens) are supplied with a spatial gradient, providing input to the gene expression system. Under these setups, the gene expression levels (protein concentrations) of each cell reach a steady state that depends on the location of the cell, leading to pattern formation[61].
Let us then assume that the fitness is given by the number of final cell types defined by the expression pattern of target genes (specifically four target genes are adopted, so that cell types are possible according to the possibilities of all their on-off patterns.) By using the fitness and mutating the gene regulation networks over generations, the individuals with higher fitness are selected in each generation. After a few hundred generations, the networks have evolved to reach the fittest. It turns out that during evolution, the expression of a certain gene (protein) slowly changes, controlling the expression of the output genes. (For details, see the original paper[62]).
@@
C) Behind the appearance of the developmental hourglass in the model is the acquisition of genes whose expression level changes slowly. Slow changes in the expression level of such genes control the timing of the developmental process and induce developmental robustness. They control the expression dynamics of other genes, which also facilitates evolution. In the model simulations, these slow expression dynamics have been conserved among organisms within a certain phylogenetic range of species sharing the hourglass (point 2).
D) Thus, the developmental hourglass can be related to robustness and dimensional reduction by a slow mode, as discussed in Β§2: The phenotypic diversity at the egg stage, derived from the mother, has to be reduced at some point in order for the developmental process to be stable (point 3). This leads to the low-dimensional constraint on the developmental trajectory of the phenotype (point 1), which results in the reduction of the variance at mid-development, the timing of which is given by the slowly changing gene expression. By assuming that is correlated with the variance over mutants through developmental time, the developmental bottleneck across species is expected (point 4)[63].
@@ 8 Collapse of dimensional reduction: Anna Karenina phenomena?
So far we have discussed the achievement of dimensional reduction in a biologically good state, in which macro-micro consistency is satisfied, i.e., cell growth and molecular replication are balanced. Then, it will be of interest to study if the reduction remains to be valid for cells with a bad, non-growing state which are put into poor environmental conditions. In fact, in bacteria, there generally exists a transition to the dormant state, i,e., the growth-arrested state, when the nutrient concentration is decreased[64, 65], whereas theoretical models for the transition from an exponential growth state to a quiescent state (which cannot grow) have also been proposed[66, 67]. Will the dimensional reduction collapse with this transition to the dormant state?
@@
If the rate of the second reaction is much faster, the product is formed as soon as the complex is formed, so that the above reaction processes are reduced to single mass action kinetics as , without intermediate complex formation. Then the model is reduced to the cell model already mentioned in Β§2-6. On the other hand, if the rate is not so fast, the intermediate complex remains for some time. These complexes can accumulate depending on the conditions, especially when the resource is limited. If such accumulation occurs, the abundances of free reactants not bound in complexes will decrease, which can hinder the reaction processes.
@@
In fact, an increase in the diversity of chemical components is observed as the cell enters a non-growing quiescent state. This increase in diversity suggests a violation of dimensional reduction in an exponential growth state. In a bad state, the dimensional reduction achieved for a robust state with consistency between macro and micro levels may collapse, leading to diversity in components. Such collapse of dimensional reduction may be also seen in the microorganism behavior[41], or possibly in cancer cells with possible increase in the phenotypic variances[69]. Probably, the first sentence of Tolstoyβs Anna Karenina, βHappy families are all alike; every unhappy family is unhappy in its own wayβ, can be applied to cells or biological systems.
@@ 9 Summary
In this paper, we have reviewed recent advances in evolutionary dimension reduction. Phenotypes in biological systems are generally robust to perturbations whereas they maintain plasticity, so that they are changed by evolution. The phenotypic changes are originally be high-dimensional, but to satisfy this robustness and plasticity, they are constrained to a low-dimensional manifold (space). In this case, the responses of phenotypes to environmental changes, to stochastic perturbations, and to genetic changes follow dominantly along this low-dimensional manifold, and thus the responses to environmental and genetic changes, as well as fluctuations to noise, and to genetic mutations are strongly correlated, allowing the prediction of phenotypic evolution. The extension of this dimensional reduction theory to the development-evolution relationship is briefly reviewed. Finally, a possible collapse of the reduction in a dormant (non-growing) cellular state is briefly discussed.